A pitcher throws a ball at 40.0 m/s and the ball is electronically timed to arrive at home plate 0.4625 s later. What is the distance from the pitcher to the home plate?
Answer #1

Solution) Given

Velocity V = 40 m/s

Time t = 0.4625 s

Distance D = ?

We know

Velocity = Distance/time

V = D/t

D = V × t

D = 40 × 0.4625

D = 18.5 m